OpenStudy (anonymous):

trying to figure out how to subtract a mixed number from a fraction like 12 2/19-17/20? Help!

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[Do you mean 12(2/19) - 17/20?\] If so convert mixed fraction into simple fraction \[12(2/19) = (2 + (12 \times 19))/19 = 230/19\] then subtract fraction 17/20 from fraction 230/19